# Upper and Lower Bounds on the Smoothed Complexity of the Simplex Method

### v2

@@ 1.2 Related work
The shadow vertex simplex algorithm has played a key role in many analyses of simplex and simplex-like algorithms. On well-conditioned polytopes, such as those of the form $\{x\in\mathbb{R}^{d}:Ax\leq b\}$ where $A$ is integral with subdeterminants bounded by $\Delta$, the shadow vertex method has been studied by in [jour/dcg/DH16, conf/icalp/BR13]. The shadow vertex method on polytopes all whose vertices are integral was studied in [simplexzeroone, black22]
On random polytopes of the form $\{x\in\mathbb{R}^{d}:Ax\leq\mathbf{1}_{n}\}$, assuming the constraint vectors are independently drawn from any rotationally symmetric distribution, the expected iteration complexity of the shadow vertex simplex method was studied by [b82, b87, b99]. In the case when the rows of $A$ arise from a Poisson distribution on the unit sphere, concentration results for the shadow size and diameter bounds were proven in [bdghl21]. The diameter of smoothed polyhedra was studied by [narayanan2021spectral], who used the shadow bound of [DH18] to show that most vertices, according to some measure, are connected by short paths.
A randomized algorithm for solving linear programs in weakly polynomial time, using the shadow vertex simplex method as a subroutine, was proposed in [conf/stoc/KS06]. The shadow vertex algorithm was recently used as part of the analysis of an interior point method by [adlnv].
@@
The construction in our lower bound is based on an adaptation of the extended formulation by [BN01] of the regular $2^{k}$-gon. They used this construction to obtain a polyhedral approximiation of the second order cone $\{x\in\mathbb{R}^{n+1}:\sum_{i=1}^{n}x_{i}^{2}\leq x_{n+1}^{2}\}$. A variant on their construction using fewer variables and inequalities was given by [glineur]. A more general construction based on reflection relations is used to construct extended formulation for the regular $2^{k}$-gon, as well as other polyhedra, in [Kaibel2013]. Extended formulations for regular $n$-gons, when $n$ is not a power of $2$, can be found in [VANDAELE2017217].
Approximations of the second order cone based on the work of [BN01, glineur] have been used to solve second order conic programs, see, e.g., [Brmann2015]. These approximations were included in the open-source MIP solver SCIP until version 7.0 [GamrathEtal2020OO, GamrathEtal2020ZR].
@@ 1.3.1 Smoothed Complexity Upper Bound
A fact used since the early days of smoothed analysis [ST04] states that the intersection polygon $Q\cap W$ is non-degenerate with probability measure $1$: every edge on $Q\cap W$ is uniquely given by the intersection between $W$ and a facet of $Q$, and every facet of $Q$ is spanned by exactly $d$ vertices. For any index set $I\in\binom{[n]}{d}$, write $E_{I}$ as the event that $\operatorname{conv}(a_{i}:i\in I)\cap W$ is an edge of $Q\cap W$. Non-degeneracy implies that every edge of $Q\cap W$ uniquely corresponds to an index set $I\in\binom{[n]}{d}$ such that $E_{I}$ holds.
Before sketching our proof, we first review the approach of [DH18], then discuss the main technical challenges in achieving an upper bound with better dependence on $\sigma$.
As a first step in [DH18], the authors replace the Gaussian distribution with a Laplace-Gaussian distribution. The latter distribution approximates the probability density of the former, in particular having nearly equivalent smoothed shadow size, while being $O(\sigma^{-1}\sqrt{d\log n})$-log-Lipschitz for any point on its domain. A probability distribution with probability density function $\mu$ is $L$-log-Lipschitz for some $L>0$ if, for any $x,y\in\mathbb{R}^{d}$, the condition $|\log(\mu(x))-\log(\mu(y))|\leq L\|x-y\|$ holds.
Next, define $\ell_{I}$ as the length of the edge on $Q\cap W$ that corresponds to $I$, i.w. when $\operatorname{conv}(a_{i}:i\in I)\cap W$ is an edge of $Q\cap W$ then $\ell_{I}$ gives the length of this line segment, and otherwise $\ell_{I}=0$. [DH18] showed that, for any family $S\subseteq\binom{[n]}{d}$, the expected number of edges of $Q\cap W$ coming from $S$ is at most
@@
where the last step comes from a Gaussian tail bound. For the denominator of (3), [DH18] showed that for any $I\in\binom{[n]}{d}$ with $\Pr[E_{I}]\geq\binom{n}{d}^{-1}$ that, conditional on $E_{I}$, the expected edge length is at least
@@
Combining the (3), (4) and (5), they obtain an upper bound of $O(\sigma^{-2}d^{2}\sqrt{\log n}+d^{3}\log^{1.5}n)$.
@@
While [DH18] made the best possible analysis based of their edge-counting strategy as outlined in (3), the strategy itself is sub-optimal. The main drawback is that using the minimum expected length of edge $\min_{I\in\binom{[n]}{d}}\mathbb{E}[\ell_{I}\mid E_{I}]$ at the denominator of (3) is too pessimistic when the edges of $Q\cap W$ are long. For instance, consider the case where an edge on $Q\cap W$ has length $\Omega(1)$ without perturbation. After the perturbation, is very likely that the length of this edge is still $\Omega(1)$, but [DH18] uses a lower-bound of $\Omega(\frac{\sigma^{2}}{d^{2}\sqrt{\log n}})$.
To improve this, we developed a new edge-counting strategy that can handle the long and short edges separately with two different ways of counting the edges. Take any index set $I\in\binom{[n]}{d}$; conditional on $E_{I}$, we write $e_{I}$ for its edge $\operatorname{conv}(a_{i}:i\in I)\cap W$. We call the next edge in clockwise direction as $e_{I^{+}}$ and say it has length $\ell_{I^{+}}$. We say $e_{I^{+}}$ is likely to be long if $\Pr[\ell_{I^{+}}\geq t\mid E_{I}]\geq 0.05$ for some parameter $t>0$ to be specified later. Let $S_{0}\subseteq\binom{[n]}{d}$ denote the set of $I\in\binom{[n]}{d}$ such that $e_{I^{+}}$ is likely to be long. Following (2), the expected number of edges in $Q\cap W$ that are likely to be long is at most
$\mathbb{E}[\sum_{I\in S_{0}}1[E_{I}]]\leq\frac{\mathbb{E}[\mathrm{perimeter}(P\cap W)]}{0.05t}\leq\frac{(2\pi+O(\sigma\sqrt{\log n}))}{0.05t}.$ (6)
Here the second step uses the exact same upper bound of $\mathbb{E}[\mathrm{perimeter}(P\cap W)]$ as in (4).
In the other case, $e_{I^{+}}$ is unlikely to be long, i.e. $\Pr[\ell_{I^{+}}\geq t\mid E_{I}]<0.05$. Now we will upper bound the number of such edges by claiming that their exterior angles each are large in expectation. Let $\theta_{I}$ be the exterior angle at the endpoint of $\operatorname{conv}(a_{i}:i\in I)\cap W$ that comes last in clockwise order, i.e., the vertex where $e_{I}$ and $e_{I^{+}}$ meet. Our key observation is that $\sin(\theta_{I})\cdot\ell_{I^{+}}$ equals the distance from the affine hull of $e_{I}$ to the second vertex of $e_{I^{+}}$ in clockwise order. Therefore, by establishing a universal lower bound for such line-to-vertex distances, we can derive a lower bound for $E[\theta_{I}\mid E_{I}]$ for any edge $e_{I^{+}}$ that is unlikely to be long. See Figure 1 for an illustration.
More formally, let $p_{I^{+}}$ denote the next vertex of $Q\cap W$ after $e_{I}$ in clockwise order. For a specific value of $\gamma>0$, Let $S_{1}\subseteq\binom{[n]}{d}$ be the collection of index sets $I\in\binom{[n]}{d}$ for which $\Pr[E_{I}]\geq 10\binom{n}{d}^{-1}$. Suppose for each $I\in S_{1}$,
@@
[…]
Figure 1: Illustration of the case when $e_{I^{+}}$ is short. In brown is the edge $e_{I}$ with its extension line dashed. The next edge in clockwise direction, drawn in black, has length $\ell_{I^{+}}$. In orange is the line-to-vertex distance $\operatorname{dist}(p_{I^{+}},\operatorname{affhull}(e_{I}))$, and in blue is the angle $\theta_{I}$. If $\operatorname{dist}(p_{I^{+}},\operatorname{affhull}(e_{I}))\geq\gamma$ then $\theta_{I}\geq\sin(\theta_{I})\geq\gamma/\ell_{I^{+}}$.
On the other hand, the sum of exterior angles of a polygon equals to $2\pi$. Therefore we can upper bound the expected number of edges that are not likely to be long by at most
@@
Summing up the number of edges induced by sets in $S_{0},S_{1}\setminus S_{0}$ and $\binom{[n]}{d}\setminus S_{1}$, we get an upper bound on the expected edge-count of $Q\cap W$ by at most
@@
where the final step follows from optimizing $t>0$ to get the strongest possible bound. We summarize our result in Theorem 28. For details of the edge-counting strategy, see Section 3.
@@
As a warm-up, we first discuss the case two-dimensional case $d=2$, which will be explained in detail in Section 4. In this case, $W$ will become the entire two-dimensional space and will disappear from consideration. Therefore, we can focus on lower-bounding the distance from any edge $e$ of the polygon $Q=\operatorname{conv}(a_{1},\ldots,a_{n})$ to any other vertices, i.e. it suffices to find $\gamma>0$ such that for any $I\in\binom{[n]}{2}$,
@@
We can obtain a lower bound on this quantity for any $L$-log-Lipschitz distribution. Through an appropriate coordinate transformation we prove that, irrespective of the values of $a_{j},j\notin I$, the distance $\operatorname{dist}(\operatorname{conv}(a_{j},j\notin I),\operatorname{affhull}(e_{I}))$, conditional on being non-zero, follows from a $2L$-log-Lipschitz distribution. We will calculate that we may choose $\gamma=\Omega(1/L)$. This result can be applied with small changes to our Gaussian random variables $a_{1},\ldots,a_{n}$ by substituting for the Laplace-Gaussian distribution following [DH18]. Plugging into (9), we get that in the two-dimensional case, $\mathbb{E}[\operatorname{edges}(Q)]\leq O(\sqrt{(1+\sigma\sqrt{\log n})/(\sigma/\sqrt{\log n})})=O(\sqrt[4]{\log n}/\sqrt{\sigma}+\sqrt{\log n})$ as stated in Theorem 4.
@@
As in the two-dimensional case, we must lower-bound of the line-to-vertex distance $\operatorname{dist}(p_{I^{+}},\operatorname{affhull}(e_{I}))$ (see (7)) of $Q\cap W$. Analyzing this, however, becomes more challenging. In two-dimensional case, each edge is the convex hull of two vertices among $a_{1},\ldots,a_{n}$ and is independent to the other vertices on $Q\cap W$. In contrast, if $d\geq 3$ then each edge on $Q\cap W$ will be the intersection between $W$ and a $(d-1)$-dimensional facet of $Q$ (which is the convex hull of $d$ vertices), and each vertex will be the intersection between $W$ and a $(d-2)$-dimensional ridge of $Q$ (which is the convex hull of $d-1$ vertices). So the distribution of $e_{I}$ and $p_{I^{+}}$ are correlated.
To overcome these difficulties, we first factor $\operatorname{dist}(p_{I^{+}},\operatorname{affhull}(e_{I}))$ into the product of separate parts which are easier to analyze, and then use log-Lipschitzness of $a_{1},\ldots,a_{n}$ to lower-bound each part with good probability. Fix without loss of generality $e=e_{[d]}=\operatorname{conv}(a_{1},\ldots,a_{d})\cap W$, as the potential edge of interest. Consider the second endpoint $p$ on $e$ in clockwise direction and let $J\in\binom{[d]}{d-1}$ be the index set such that $\{p\}=\operatorname{conv}(a_{j}:j\in J)\cap W$. Let $p^{\prime}=\operatorname{conv}(a_{i}:i\in J^{\prime})\cap W$ (with $J^{\prime}\in\binom{[n]}{d-1}$) be the vertex next to the edge $e$ in clockwise direction. From the non-degeneracy conditions, we know that $J^{\prime}$ only differs to $J$ with two vertices almost surely, so we can assume without loss of generality that $J=\{2,\ldots,{d}\}$ and $J^{\prime}=\{3,\ldots,d\}\cup\{k\}$ for some $k\in\{d+1,\ldots,n\}$.
The main idea of our analysis is the observation that if the radius of $Q$ is bounded above by $O(1)$ (which happens with overwhelming probability due to Gaussian tail bound), then we can lower bound the two-dimensional line-to-vertex distance $\operatorname{dist}(p^{\prime},\operatorname{affhull}(e))$ by the product of two distances $\Omega(\delta\cdot r)$, where
  • $\delta$ is the $d$-dimensional distance from the facet-defining hyperplane $\operatorname{affhull}(a_{1},\ldots,a_{d})$ containing $e$, to the vertices that are not in the facet, i.e.
@@
  • $r$ is the distance from the boundary of the ridge $\partial\operatorname{conv}(a_{2},\ldots,a_{d})$ to the one-dimensional line $\operatorname{affhull}(e)$, i.e. $r=\operatorname{dist}(\operatorname{affhull}(e),\partial\operatorname{conv}(a_{2},\ldots,a_{d}))$.
We will give the formal statement of the distance splitting lemma in Lemma 36.
@@
  • The lower bound of $r$ resembles a more technical version of the proof of the “distance lemma” of [ST04].
@@
    First we show that each vertex of the ridge $\operatorname{conv}(\pi(a_{2}),\ldots,\pi(a_{d}))$ is $\Omega(1/(d^{2}L))$-far away from the hyperplane spanned by its other vertices. That means that the projected ridge $\operatorname{conv}(\pi(a_{2}),\dots,\pi(a_{d}))$ is wide in every direction.
@@
We conclude our main result of the line-to-vertex distance lower bound in Lemma 33. Readers are referred to Section 5 for detailed discussions.

### v3

@@ 1.2 Related work
The shadow vertex simplex algorithm has played a key role in many analyses of simplex and simplex-like algorithms. On well-conditioned polytopes, such as those of the form $\{x\in\mathbb{R}^{d}:Ax\leq b\}$ where $A$ is integral with subdeterminants bounded by $\Delta$, the shadow vertex method has been studied by in [15, 16]. The shadow vertex method on polytopes all whose vertices are integral was studied in [9, 10]
On random polytopes of the form $\{x\in\mathbb{R}^{d}:Ax\leq\mathbf{1}_{n}\}$, assuming the constraint vectors are independently drawn from any rotationally symmetric distribution, the expected iteration complexity of the shadow vertex simplex method was studied by [14, 13, 12]. In the case when the rows of $A$ arise from a Poisson distribution on the unit sphere, concentration results for the shadow size and diameter bounds were proven in [11]. The diameter of smoothed polyhedra was studied by [44], who used the shadow bound of [17] to show that most vertices, according to some measure, are connected by short paths.
A randomized algorithm for solving linear programs in weakly polynomial time, using the shadow vertex simplex method as a subroutine, was proposed in [39]. The shadow vertex algorithm was recently used as part of the analysis of an interior point method by [3].
@@
The construction in our lower bound is based on an adaptation of the extended formulation by [8] of the regular $2^{k}$-gon. They used this construction to obtain a polyhedral approximation of the second order cone $\{x\in\mathbb{R}^{n+1}:\sum_{i=1}^{n}x_{i}^{2}\leq x_{n+1}^{2}\}$. A variant on their construction using fewer variables and inequalities was given by [28]. A more general construction based on reflection relations is used to construct extended formulation for the regular $2^{k}$-gon, as well as other polyhedra, in [37]. Extended formulations for regular $n$-gons, when $n$ is not a power of $2$, can be found in [51].
Approximations of the second order cone based on the work of [8, 28] have been used to solve second order conic programs, see, e.g., [7]. These approximations were included in the open-source MIP solver SCIP until version 7.0 [26].
@@ 1.3.1 Smoothed Complexity Upper Bound
A fact used since the early days of smoothed analysis [49] states that the intersection polygon $Q\cap W$ is non-degenerate with probability measure $1$: every edge on $Q\cap W$ is uniquely given by the intersection between $W$ and a facet of $Q$, and every facet of $Q$ is spanned by exactly $d$ vertices. For any index set $I\in\binom{[n]}{d}$, write $E_{I}$ as the event that $\operatorname{conv}(a_{i}:i\in I)\cap W$ is an edge of $Q\cap W$. Non-degeneracy implies that every edge of $Q\cap W$ uniquely corresponds to an index set $I\in\binom{[n]}{d}$ such that $E_{I}$ holds.
Before sketching our proof, we first review the approach of [17], then discuss the main technical challenges in achieving an upper bound with better dependence on $\sigma$.
As a first step in [17], the authors replace the Gaussian distribution with a distribution they call the Laplace-Gaussian distribution [^2] The Laplace-Gaussian distribution follows the Gaussian probability density function near its mean but exhibits exponentially decaying tails.. The latter distribution approximates the probability density of the former, in particular having nearly equivalent smoothed shadow size, while being $O(\sigma^{-1}\sqrt{d\log n})$-log-Lipschitz for any point on its domain. A probability distribution with probability density function $\mu$ is $L$-log-Lipschitz for some $L>0$ if, for any $x,y\in\mathbb{R}^{d}$, the condition $|\log(\mu(x))-\log(\mu(y))|\leq L\|x-y\|$ holds.
Next, define $\ell_{I}$ as the length of the edge on $Q\cap W$ that corresponds to $I$, i.e., when $\operatorname{conv}(a_{i}:i\in I)\cap W$ is an edge of $Q\cap W$ then $\ell_{I}$ gives the length of this line segment, and otherwise $\ell_{I}=0$. [17] showed that, for any family $S\subseteq\binom{[n]}{d}$, the expected number of edges of $Q\cap W$ coming from $S$ is at most
@@
where the last step comes from a Gaussian tail bound. For the denominator of (3), [17] showed that for any $I\in\binom{[n]}{d}$ with $\Pr[E_{I}]\geq\binom{n}{d}^{-1}$ that, conditional on $E_{I}$, the expected edge length is at least
@@
By combining (3), (4) and (5), they proved an upper bound of $O(\sigma^{-2}d^{2}\sqrt{\log n}+d^{3}\log^{1.5}n)$.
@@
While [17] made the best possible analysis based of their edge-counting strategy as outlined in (3), the strategy itself is sub-optimal. The main drawback is that using the minimum expected length of edge $\min_{I\in\binom{[n]}{d}}\mathbb{E}[\ell_{I}\mid E_{I}]$ at the denominator of (3) is too pessimistic when the edges of $Q\cap W$ are long. For instance, consider the case where an edge on $Q\cap W$ has length $\Omega(1)$ without perturbation. After the perturbation, is very likely that the length of this edge is still $\Omega(1)$, but [17] uses a lower-bound of $\Omega(\frac{\sigma^{2}}{d^{2}\sqrt{\log n}})$.
To improve this, we developed a new edge-counting strategy that can handle the long and short edges separately with two different ways of counting the edges. Take any index set $I\in\binom{[n]}{d}$; conditional on $E_{I}$, we write $e_{I}$ for its edge $\operatorname{conv}(a_{i}:i\in I)\cap W$. We call the next edge in clockwise direction as $e_{I^{+}}$ and say it has length $\ell_{I^{+}}$. We say $e_{I^{+}}$ is likely to be long if $\Pr[\ell_{I^{+}}\geq t\mid E_{I}]\geq 0.05$ for some parameter $t>0$ to be specified later. Here $0.05$ can be replaced by some arbitrary constant in $(0,0.5)$. Let $S_{0}\subseteq\binom{[n]}{d}$ denote the set of $I\in\binom{[n]}{d}$ such that $e_{I^{+}}$ is likely to be long. Following (2), the expected number of edges in $Q\cap W$ that are likely to be long is at most
$\mathbb{E}[\sum_{I\in S_{0}}1[E_{I}]]\leq\frac{\mathbb{E}[\mathrm{perimeter}(Q\cap W)]}{0.05t}\leq\frac{(2\pi+O(\sigma\sqrt{\log n}))}{0.05t}.$ (6)
Here the second step uses the exact same upper bound of $\mathbb{E}[\mathrm{perimeter}(Q\cap W)]$ as in (4).
In the other case, $e_{I^{+}}$ is unlikely to be long, i.e., $\Pr[\ell_{I^{+}}\geq t\mid E_{I}]<0.05$. Now we will upper bound the number of such edges by claiming that their exterior angles each are large in expectation. Let $\theta_{I}$ be the exterior angle at the endpoint of $\operatorname{conv}(a_{i}:i\in I)\cap W$ that comes last in clockwise order, i.e., the vertex where $e_{I}$ and $e_{I^{+}}$ meet. Our key observation is that $\sin(\theta_{I})\cdot\ell_{I^{+}}$ equals the distance from the affine hull of $e_{I}$ to the second vertex of $e_{I^{+}}$ in clockwise order. Therefore, by establishing a universal lower bound for such line-to-vertex distances, we can derive a lower bound for $E[\theta_{I}\mid E_{I}]$ for any edge $e_{I^{+}}$ that is unlikely to be long. See Figure 1 for an illustration.
More formally, let $p_{I^{+}}$ denote the next vertex of $Q\cap W$ after $e_{I}$ in clockwise order.
Let $S_{1}\subseteq\binom{[n]}{d}$ be the collection of index sets $I\in\binom{[n]}{d}$ for which $\Pr[E_{I}]\geq 10\binom{n}{d}^{-1}$. For a specific value of $\gamma>0$, suppose that for each $I\in S_{1}$ we have
@@
Figure 1: Illustration of the case when $e_{I^{+}}$ is short. In purple is the edge $e_{I}$ with its extension line dashed. The next edge in clockwise direction, $e_{I}^{+}$, has length $\ell_{I^{+}}$ and is drawn in black. In red is the line-to-vertex distance $\operatorname{dist}(p_{I^{+}},\operatorname{affhull}(e_{I}))$, and in blue is the angle $\theta_{I}$. If $\operatorname{dist}(p_{I^{+}},\operatorname{affhull}(e_{I}))\geq\gamma$ then $\theta_{I}\geq\sin(\theta_{I})\geq\gamma/\ell_{I^{+}}$.
On the other hand, the sum of exterior angles of a polygon equals to $2\pi$. Therefore we can upper bound the expected number of edges that are not likely to be long of at most
@@
Summing up the number of edges induced by sets in $S_{0},S_{1}\setminus S_{0}$ and $\binom{[n]}{d}\setminus S_{1}$, we get an upper bound on the expected edge-count of $Q\cap W$ of at most
@@
where the final step follows from optimizing $t>0$ to get the strongest possible bound. We summarize our result in Theorem 3.4. For details of the edge-counting strategy, see Section 3.
@@
As a warm-up, we first discuss the two-dimensional case $d=2$, which will be explained in detail in Section 4. In this case, $W$ will become the entire two-dimensional space and will disappear from consideration. Therefore, we can focus on lower-bounding the distance from any edge $e$ of the polygon $Q=\operatorname{conv}(a_{1},\ldots,a_{n})$ to any other vertices, i.e., it suffices to find $\gamma>0$ such that for any $I\in\binom{[n]}{2}$,
@@
We can obtain a lower bound on this quantity for any $L$-log-Lipschitz distribution. Through an appropriate coordinate transformation we prove that, irrespective of the values of $a_{j},j\notin I$, the distance $\operatorname{dist}(\operatorname{conv}(a_{j},j\notin I),\operatorname{affhull}(e_{I}))$, conditional on being non-zero, follows from a $2L$-log-Lipschitz distribution. We will calculate that we may choose $\gamma=\Omega(1/L)$. This result can be applied with small changes to our Gaussian random variables $a_{1},\ldots,a_{n}$ by substituting for the Laplace-Gaussian distribution following [17]. Plugging into (9), we get that in the two-dimensional case, $\mathbb{E}[\operatorname{edges}(Q)]\leq O(\sqrt{(1+\sigma\sqrt{\log n})/(\sigma/\sqrt{\log n})})=O(\sqrt[4]{\log n}/\sqrt{\sigma}+\sqrt{\log n})$ as stated in Theorem 1.4.
@@
As in the two-dimensional case, we must lower-bound of the line-to-vertex distance $\operatorname{dist}(p_{I^{+}},\operatorname{affhull}(e_{I}))$ (see (7)) of $Q\cap W$. Analysing this, however, becomes more challenging. In two-dimensional case, each edge is the convex hull of two vertices among $a_{1},\ldots,a_{n}$ and is independent of the other potential vertices $a_{i}$. In contrast, if $d\geq 3$ then each edge on $Q\cap W$ will be the intersection between $W$ and a (d1)(d-1)-dimensional facet of $Q$ (which is the convex hull of $d$ vertices), and each vertex will be the intersection between $W$ and a (d2)(d-2)-dimensional ridge of $Q$ (which is the convex hull of d1d-1 vertices). So the distributions of $e_{I}$ and $p_{I^{+}}$ are correlated.
[…]
The main idea of our analysis is the observation that if the Euclidean diameter of $Q$ is bounded above by $O(1)$ (which happens with overwhelming probability due to Gaussian tail bound), then we can lower bound the two-dimensional line-to-vertex distance $\operatorname{dist}(p^{\prime},\operatorname{affhull}(e))$ by the product of two distances $\Omega(\delta\cdot r)$, where
  • $\delta$ is the $d$-dimensional distance from the facet-defining hyperplane $\operatorname{affhull}(a_{1},\ldots,a_{d})$ containing $e$, to the vertices of $Q$ that are not in the facet, i.e.,
@@
  • $r$ is the distance from the boundary of the ridge $\partial\operatorname{conv}(a_{2},\ldots,a_{d})$ to the one-dimensional line $\operatorname{affhull}(e)$, i.e., $r=\operatorname{dist}(\operatorname{affhull}(e),\partial\operatorname{conv}(a_{2},\ldots,a_{d}))$.
We will give the formal statement of the distance splitting lemma in Lemma 5.6.
@@
  • The lower bound of $r$ resembles a more technical version of the proof of the “distance lemma” of [49].
@@
    First we show that each vertex of the projected ridge $\operatorname{conv}(\pi(a_{2}),\ldots,\pi(a_{d}))$ is a distance $\Omega(1/(d^{2}L))$ away from the hyperplane spanned by its other vertices. That means that the projected (d2)(d-2)-dimensional ridge $\operatorname{conv}(\pi(a_{2}),\dots,\pi(a_{d}))$ is likely to be wide in every direction.
@@
We conclude our main result of the line-to-vertex distance lower bound in Lemma 5.2. Readers are referred to Section 5 for detailed discussions.